show that the straight line (a+2b)x+(a-3b)y+b-a=0 always passes through a fixed point: find the coordinate of that point

Grade:11

1 Answers

Arun
5 years ago
 
Dear student
(x+y-1)a + (2x-3y+1) b = 0
Hence
x +y-1 = 0
x +y = 1  and  2x -3y = -1
On solving
5x = 2
x = 2/5
and 
y = 3/5